OpenStudy (anonymous):

BEN IS MAKING WAFFLES FOR HIS 3 FRIENDS .THE RECIPE CALLS FOR 2/3 CUP OF BLEACHED FLOUR AND 1/4 CUP OF WHEAT FLOUR .HOW MANY CUPS OF BLEACHED FLOUR DOES BEEN NEED ?

OpenStudy (anonymous):

he only needs 2/3 c of bleach flour. I think your question is possibly not correct

OpenStudy (anonymous):

you have not mentioned that the given recipe is for how many people.. if it is for one. then ben needs 2 cups of bleach flour

OpenStudy (anonymous):

i have seen this question before - i think what she is meaning to ask is how much total flour does he need for the recipe 2/3 + 1/4 = 8/12 + 3/12 = 11/12